What volume of hydrogen at s.t.p will be produced when 48g of magnesium reacts completely with dilute hydroxhloric according to the following equation Mg+2HCL=MgCl+H
1
Expert's answer
2020-02-25T05:02:50-0500

According to the following equation Mg+2HCL=MgCl2+H2 1 mole of magnesium produces 1 mole of hidrogen. From an equation n=m/Mr we can find the quantity of Mg. n(Mg)= 48/24=2 mole. So 2 mole of Mg produces 2 mole of H2. 1 mole of gas at s.t.p has a volume of 22.4 liters, so volume of 2 mole H2 is 2*22.4=44.8 liters.

n=m/Mr;

n=V/Vm; Vm=22.4 liters

V/Vm = m/Mr;

V=m*Vm/Mr;

V=48*22.4/24 = 44.8 l